OpenStudy (anonymous):

Determine the prime factorization of 90. A. 2 • 3^2 • 5 B. 2^3 • 11 C. 2^2 • 3 • 5 D. 2^2 • 11

OpenStudy (anonymous):

a

OpenStudy (zzr0ck3r):

I always start by trying to cut it in half, if I can 90 = 45*2 now 2 is prime so we are done there, but 45 = 9*5 and 5 is prime so we are done there, so we look at 9 = 3*3 so 90=45*2 = 9*5*2 = 3*3*5*2 = 3^2*5*2

OpenStudy (zzr0ck3r):

@song6 read the rules, that does not help anyone....

OpenStudy (anonymous):

thank you. I got messed up since I thought 3^2 was 6, those always confuse me

OpenStudy (zzr0ck3r):

easy thing to do...

OpenStudy (anonymous):

ok....